OpenStudy (anonymous):

what is e^j5t(1+j sqrt(3)) ?

OpenStudy (anonymous):

when i turn (1+j sqrt(3)) into polar from is gives me 2(cos2pi3 +sin2pi/3)

OpenStudy (anonymous):

@ganeshie8 could you please explain this?

OpenStudy (misty1212):

\[ e^{j5t(1+j \sqrt(3)} \] confused

OpenStudy (misty1212):

\[1+i\sqrt3=2(\cos(\frac{\pi}{3})+i\sin(\frac{\pi}{3}))\] is correct though

OpenStudy (misty1212):

or if you prefer \[\huge 2e^{\frac{\pi}{3}i}\]

OpenStudy (misty1212):

is the first part \[e^{5jt}\]? what is \(t\)?

OpenStudy (anonymous):

that what i got, but apparently the answer is

OpenStudy (phi):

they made a mistake. \( 1 + j \sqrt{3} = 2 e^{j \frac{\pi}{3}} \) not \( 2 e^{j \frac{\pi}{6}} \)